# reference.bib

### v1

@@ 8.1. Examples with rank 11 11
    But for $p\neq 3$, $\sqrt{-3}$ is a unit in $\mathcal{O}_{K,p}$. So, the condition becomes $x-1\in 2\mathcal{O}_{K,p}\textup{ }\forall p\neq 3$. We can then check that this is true only for ±$\pm 1\in U$. Therefore, $[U:U^{\prime}]=3$. Also, $\mu=1$ in this case, because 33 33 ramifies in $K$. Also, as $-3\equiv 5\pmod{8}$, we have that 22 22 remains prime in $K$, hence
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=(1/1)\cdot 2^{1-1-1}\cdot 2\cdot 3^{-1}\cdot 2\cdot(1+1/2)=1.$
  • Assume $t>3$ with $-t\nequiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ prime factors. In this case, $\mathfrak{e}=\mathfrak{c}=\mathbb{Z}$ and $\mathfrak{a}=t\mathbb{Z}$. For all $p_{i}$, we have that $r_{p_{i}}(\xi)$ satisfies the conditions of [shimura_diophantine_2006, Theorem 5.7] and therefore $\mathfrak{a}^{*}=t\mathbb{Z}$. We have $\mu=k$ as each $p_{i}$ is ramified in $K$. Now $U=\mathcal{O}_{K}^{\times}=\{\pm 1\}$ and
@@
    But for all $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore ±$\pm 1\in U^{\prime}$, i.e. $[U:U^{\prime}]=1$. We then obtain:
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=c_{K}\cdot 2^{1-k-1}\cdot 2=c_{K}\cdot 2^{1-k}$
    and therefore this is 11 11 if $c_{K}=2^{k-1}$. Hence, the answer in this case is the number fields $K=\mathbb{Q}(\sqrt{-t})$ that satisfy $c_{K}=2^{k-1}$, where $k$ is the number of prime factors of $t$ and $-t\nequiv 1\pmod{4}$. For example, when $c_{K}=2$, $t$ must have 2 prime factors, and examples would be $t=6,10,$ etc.
  • Finally, assume $t>3$ with $-t\equiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ primes. In these cases, similarly to the case $t=3$, we have $\mathfrak{a}=t\mathbb{Z}$, $\mathfrak{a}^{*}=t\mathbb{Z}$ and $\mathfrak{c}=2\mathbb{Z}$. Also, $\mu=k$ because each prime $p_{i}$ is ramified in $K$. Now, $U=\mathcal{O}_{K}^{\times}=\{\pm{1}\}$ and then
@@
    But for $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore ±$\pm 1\in U^{\prime}$, which means $[U:U^{\prime}]=1$. Now, if $-t\equiv 5\pmod{8}$, we get that 22 22 is inert in $K$ and if $-t\equiv 1\pmod{8}$, 22 22 splits in $K$. Therefore
    $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=\begin{cases}c_{K}\cdot 2^{1-k-1}\cdot 2\cdot 2\cdot(1+1/2)=3c_{K}\cdot 2^{1-k}&\textup{ if }-t\equiv 5\pmod{8}\\ c_{K}\cdot 2^{1-k-1}\cdot 2\cdot 2\cdot(1-1/2)=c_{K}\cdot 2^{1-k}&\textup{ if }-t\equiv 1\pmod{8}.\end{cases}$
    In the first case the index cannot be 11 11, while in the second case, we must have $c_{K}=2^{k-1}$. Therefore, the answer in this case is $t$ such that $-t\equiv 1\pmod{8}$ so that if $K=\mathbb{Q}(\sqrt{-t})$ we have $c_{K}=2^{k-1}$, where $k$ is the number of distinct prime factors of $t$. For example, the only example for $k=2$ is $t=15$.
@@
Let now $t$ be one of the above. We aim to give an Euler product expression for $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}$. In particular, from Theorem 7.2, we need to give an Euler product expression for
$\zeta_{\xi,\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi;N)N^{-s},$
where now $n(\xi;N)=\#\{s\in\mathbb{Z}/2tN\mathbb{Z}\mid s^{2}\equiv-4t\pmod{4tN}\}$, as we have explained in Remark 7.
@@
Now, for all $p\notin\mathcal{P}$, we have $(p,t)=1$, so by [congruences, Proposition 14], we get that
@@
for all $k\geq 1$, where $\left(\frac{\cdot}{p}\right)$ denotes the Legendre symbol. Therefore, if we define $\chi_{t}(p):=\left(\frac{-t}{p}\right)$, we deduce that (bear in mind that $\chi_{t}^{2}=1$):
$\zeta_{\xi,\mathcal{P}}(s)=\zeta_{\mathcal{P}}(s)\zeta_{\mathcal{P}}(2s)^{-1}\zeta_{\mathcal{P}}(s,\chi_{t}).$
Therefore, we have the following theorem:
@@
Let $S=2t$, with $t$ being chosen as in Proposition 8.1. We define $\chi_{t}(p):=\left(\frac{-t}{p}\right)$ and $\psi(p):=\left(\frac{-1}{p}\right)$. Let also $\mathcal{P}$ a sufficiently large finite set of primes, containing the primes described in Theorem 7.2 and the primes so that the conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. We then have:
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}=A(\xi)L_{\mathcal{P}}\left(F;s-k+\frac{3}{2}\right)\zeta_{\mathcal{P}}(2s-2k+4)^{-1}\frac{\zeta_{\mathcal{P}}(s-k+2,\chi_{t})}{\zeta_{\mathcal{P}}(s-k+2,\psi)},$
where again $\xi=(1,0,1)^{t}$.
@@
The proof follows by Theorem 7.2 after choosing $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. We have computed $\zeta_{\xi,\mathcal{P}}$ above and
@@
as this can be computed by [shimura_orthogonal, Proposition 5.15], because $\mathcal{P}$ is chosen so that conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. ∎
@@
In the case $t=1$, we recover the result of Kohnen and Skoruppa in [kohnen_skoruppa]. For $t>1$, the result can be interpreted as a result of modular forms for a paramodular group (cf. [krieg_integral_orthogonal, Corollary 6], [paramodular_hecke]).
@@ 8.2. The rank $n\geq 2$ case
In the rank $n\geq 2$ case, a Theorem like [shimura_diophantine_2006, Theorem 5.7] is not available. For this reason, we seek examples of matrices $S$ so that the following conditions hold:
@@
    The number of classes in the genus of maximal lattices (this is independent of the choice of the maximal lattice, see [shimura2004arithmetic, Paragraph 9.7]) in $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$ is 11 11. Here, $\psi:=\phi_{0}\mid_{W}$.
We will show that for rank $n=2,4,6,8$, there is at least one positive definite even symmetric matrix $S$ of rank $n$, so that the above conditions are satisfied. The reasons why we only provide these examples is explained in Section 9. We start with the following lemma:
@@
We have $H(\xi)_{\mathbb{Q}}=SO^{\psi}(W)$ with $\psi$ the bilinear form induced by the matrix:
$T=\frac{1}{2}\begin{pmatrix}-2&\\ &-S\end{pmatrix}.$
@@
We have $W=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$. Then, it is easy to see that
$W=\textup{span}\{(1,\textbf{0},-1),(0,1,0,\cdots,0,0),(0,0,1,\cdots,0,0),\cdots,(0,0,0,\cdots,1,0)\}.$
By then evaluating $\phi_{0}[x]$ for the above basis vectors $x$ of $W$, the lemma follows. ∎
@@
  • $n=6$: $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$ with determinant 33 33.
  • $n=8$: $S=\begin{pmatrix}2&-1&1&1&-1&-1&1&-1\\ -1&2&0&-1&0&1&-1&1\\ 1&0&2&1&-1&0&0&0\\ 1&-1&1&2&-1&-1&0&-1\\ -1&0&-1&-1&2&1&-1&1\\ -1&1&0&-1&1&2&-1&1\\ 1&-1&0&0&-1&-1&2&-1\\ -1&1&0&-1&1&1&-1&2\end{pmatrix}$ with determinant 11 11.
@@
The matrices $S$ of rank 22 22 correspond to the unitary groups of the imaginary quadratic fields $\mathbb{Q}(\sqrt{-3}),\mathbb{Q}(\sqrt{-15})$ respectively (cf. [eisenstein_thesis, Example 1.6.6, (v)]).
Let us first check the conditions $(1)$ and $(3)$. Condition $(1)$ follows by [eisenstein_thesis, Proposition 1.6.12] for the matrices with square-free determinant and [eisenstein_thesis, Lemma 1.6.5] for the matrix of determinant 99 99. Condition $(3)$ follows by [hanke2011enumerating, Section 8], as $W$ corresponds to the definite orthogonal group induced by the matrices:
  • $n=2$: Examples $4,26$ in matrices of 33 33 variables.
  • $n=4$: Examples $3,5$ in matrices of 55 55 variables.
  • $n=6$: Example 33 33 in matrices of 77 77 variables.
  • $n=8$: Example 11 11 in matrices of 99 99 variables.
All these examples correspond to one class in the genus of the standard lattice $\mathbb{Z}^{n+1}$, in the cases when it is maximal. This can also be seen by a simple computation in MAGMA for example. As we mentioned above, this shows that every maximal lattice has one class in its genus. Finally, we need to check condition $(2)$. We will use [shimura2004arithmetic, Proposition 11.12]. We have the following local result:
@@
Our proof is based on [shimura2004arithmetic, Proposition 11.12]. We will first establish the following claim:
@@
Indeed, assume $p\nmid\det(S)$. Then, as $L_{0,p}\subset L_{0,p}^{*}$ always, suffices to show that $L_{0,p}^{*}\subset L_{0,p}$ too. But $L_{0,p}^{*}=S_{0}^{-1}L_{0,p}$ and so if $y\in L_{0,p}^{*}$, we write $y=S_{0}^{-1}x$ with $x\in L_{0,p}$ and so the result follows as $p\nmid\det(S)$ (we have $S_{0}^{-1}=\begin{pmatrix}&&1\\ &-S^{-1}&\\ 1&&&\end{pmatrix}$).
On the other hand, if $p\mid\det(S)$, we can always find $y\in L_{0,p}^{*}\backslash L_{0,p}$.
We remind ourselves that $\phi_{0}[\xi]=1$ and $\phi_{0}(\xi,L_{0,p})=\frac{1}{2}\mathbb{Z}_{p}$ for all primes $p$. Therefore, $\phi_{0}(\xi,L_{0,p})^{2}=\phi_{0}[\xi]\mathbb{Z}_{p}$ for all $p\neq 2$. Moreover, $L_{0,2}^{*}=L_{0,2}$ for every choice of $S$, as $2\nmid\det(S)$ for any $S$. This means [shimura2004arithmetic, Proposition 11.12] is applicable in every case.
Let now $t_{p}$ denote the dimension of the maximal anisotropic subspace of $(\mathbb{Q}_{p},\phi_{0})$ (cf. [shimura2004arithmetic, Paragraph 8.3]). For $p\nmid\det(S)$, we have $L_{0,p}^{*}=L_{0,p}$, $t_{p}\neq 1$ as $n$ is even and $4\phi_{0}[\xi]^{-1}\phi_{0}(\xi,L_{0,p})^{2}=\mathbb{Z}_{p}$. So, by [shimura2004arithmetic, Proposition 11.12, (iii), (2)], $L_{0,p}\cap W_{p}$ is $\mathbb{Z}_{p}-$maximal.
If now $p\mid\det(S)$, we claim $t_{p}>1$ and therefore [shimura2004arithmetic, Proposition 11.12, (iii), (1)] will be applicable. We define $K_{0}:=\mathbb{Q}_{p}(\sqrt{\delta})$, where $\delta:=(-1)^{(n+2)(n+1)/2}\det(\phi_{0})$. By the proof of [shimura_classification_quadratic_forms, Lemma 3.3], we have $t_{p}=2$ if and only if $K_{0}\neq\mathbb{Q}_{p}$. We show this case by case:
  • $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. Then, we claim that $t_{3}=2$. Now, $\det(\phi_{0})=-\frac{3}{2^{4}}$ and then $K_{0}=\mathbb{Q}_{3}(\sqrt{\det(\phi_{0})})=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$.
  • $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{15}{2^{4}}$ and we claim $t_{3}=t_{5}=2$. But again, if $p\in\{3,5\}$, we have $K_{0}=\mathbb{Q}_{p}(\sqrt{\det{\phi_{0}}})=\mathbb{Q}_{p}(\sqrt{-15})\neq\mathbb{Q}_{p}$ by taking valuations (for example, if $\sqrt{-15}\in\mathbb{Q}_{3}$, then $-15=u^{2}$ for some $u\in\mathbb{Q}_{3}$ and so $1\displaystyle 2v_{3}(u)=$ 1, contradiction).
  • $S=\begin{pmatrix}2&-1&-1&-1\\ -1&2&1&0\\ -1&1&2&0\\ -1&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{5}{2^{6}}$ and so $K_{0}=\mathbb{Q}_{5}(\sqrt{5})\neq\mathbb{Q}_{5}$, so $t_{5}=2$.
  • $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, $K_{0}=\mathbb{Q}_{3}$ and we can compute $t_{3}=4$ in Sage.
  • $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-\frac{3}{2^{8}}$ and so again $K_{0}=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$. Therefore, $t_{3}=2$.
@@
Finally, the fact that $D_{p}=H(\xi)_{p}\cap C_{p}$ follows from [shimura2004arithmetic, Proposition 11.12, (iv)], as $t_{p}\neq 1$ always, because $n$ is even. ∎
@@
The first claim follows by [eisenstein_thesis, Proposition 1.6.9], as all the localizations are maximal by Proposition 8.2. For the second one, we have
@@
But the lattice $xL_{0}$ is the lattice which is defined by $(xL_{0})_{p}=x_{p}L_{0,p}$ for all primes $p$. Now, if $x\in H(\xi)_{\mathbb{A}}$ with $xL_{0}=L_{0}$, we have $x(L_{0}\cap W)=L_{0}\cap W$ (cf. [shimura2004arithmetic, page 104]), so $H(\xi)_{\mathbb{A}}\cap C\subset D$.
On the other hand, if $x\in D$, then $x_{p}\in D_{p}=H(\xi)_{p}\cap C_{p}$ by Proposition 8.2 and therefore $x_{p}L_{0,p}=L_{0,p}$ for all primes $p$. This means that $x\in H(\xi)_{\mathbb{A}}\cap C$, as wanted. ∎

### v2

@@ 9.1. Examples with rank $1$
    But for $p\neq 3$, $\sqrt{-3}$ is a unit in $\mathcal{O}_{K,p}$. So, the condition becomes $x-1\in 2\mathcal{O}_{K,p}\textup{ }\forall p\neq 3$. We can then check that this is true only for $\pm 1\in U$. Therefore, $[U:U^{\prime}]=3$. Also, $\mu=1$ in this case, because $3$ ramifies in $K$. Also, as $-3\equiv 5\pmod{8}$, we have that $2$ remains prime in $K$, hence
    $H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=1\cdot 2^{1-1}\cdot 3^{-1}\cdot 2\cdot(1+1/2)=1.$
  • $t=2$ or $t>3$ with $-t\nequiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ distinct prime factors. In this case, $\mathfrak{e}=\mathfrak{c}=\mathbb{Z}$ and $\mathfrak{a}=t\mathbb{Z}$. For all $p_{i}$, we have that $r_{p_{i}}(\xi)$ satisfies the conditions of Theorem 9.1 and therefore $\mathfrak{a}^{*}=t\mathbb{Z}$. We have $\mu=k$ as each $p_{i}$ is ramified in $K$. Now $U=\mathcal{O}_{K}^{\times}=\{\pm 1\}$ and
@@
    But for all $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore $\pm 1\in U^{\prime}$, i.e. $[U:U^{\prime}]=1$. We then obtain:
    $H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]=c_{K}\cdot 2^{1-k}.$
    Therefore, this is $1$ iff $c_{K}=2^{k-1}$. Hence, the answer in this case is the number fields $K=\mathbb{Q}(\sqrt{-t})$ that satisfy $c_{K}=2^{k-1}$, where $k$ is the number of prime factors of $t$ and $-t\nequiv 1\pmod{4}$. For example, when $c_{K}=2$, $t$ must have 2 prime factors, and examples would be $t=2,6,10,$ etc.
  • $t>3$ with $-t\equiv 1\pmod{4}$. We write $t=p_{1}\cdots p_{k}$ with $p_{i}$ primes. In these cases, similarly to the case $t=3$, we have $\mathfrak{a}=t\mathbb{Z}$, $\mathfrak{c}=2\mathbb{Z}$ and $\mathfrak{a}^{*}=t\mathbb{Z}$. Also, $\mu=k$ because each prime $p_{i}$ is ramified in $K$. Now, $U=\mathcal{O}_{K}^{\times}=\{\pm{1}\}$ and then
@@
    But for $p\neq p_{i}$, $\sqrt{-t}$ is a unit in $\mathcal{O}_{K,p}$ and therefore $\pm 1\in U^{\prime}$, which means $[U:U^{\prime}]=1$. Now, if $-t\equiv 5\pmod{8}$, we get that $2$ is inert in $K$ and if $-t\equiv 1\pmod{8}$, $2$ splits in $K$. Therefore
[…]
    In the first case the index cannot be $1$, while in the second case, we must have $c_{K}=2^{k-1}$. Therefore, the answer in this case is $t$ such that $-t\equiv 1\pmod{8}$ so that if $K=\mathbb{Q}(\sqrt{-t})$ we have $c_{K}=2^{k-1}$, where $k$ is the number of distinct prime factors of $t$. For example, the only example for $k=2$ is $t=15$.
@@
Let now $t$ be one of the above. The set of primes $\mathcal{P}$ is as described in Theorem 8.2 and we include the prime $2$. We aim to give an Euler product expression for $\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)$. In particular, from Theorem 8.2, we need to give an Euler product expression for
where now $n(\xi;N)=\#\{s\in\mathbb{Z}/2tN\mathbb{Z}\mid s^{2}\equiv-4t\pmod{4tN}\}$, as we explained in the beginning of Section 9.
@@
Now, for all $p\notin\mathcal{P}$, we have $(p,t)=1$, so from [29, Proposition 14] (as we assume $2\in\mathcal{P}$), we get
@@
for all $k\geq 1$, where $\left(\frac{\cdot}{p}\right)$ denotes the Legendre symbol. Therefore, if we define $\chi_{t}(p):=\left(\frac{-t}{p}\right)$ for $p\not\in\mathcal{P}$, we deduce that (bear in mind that $\chi_{t}^{2}=1$)
$\zeta_{\xi,\mathcal{P}}(s)=\zeta_{\mathcal{P}}(s)\zeta_{\mathcal{P}}(2s)^{-1}\zeta_{\mathcal{P}}(s,\chi_{t}),$ (20)
where $\zeta_{\mathcal{P}}(s,\chi_{t}):=\prod_{p\not\in\mathcal{P}}(1-\chi_{t}(p)p^{-s})^{-1}$. Therefore, we have the following Theorem.
@@
Let $S=2t$, with $t$ being chosen as in Proposition 9.1. Assume $\mathcal{P}$ is a finite set of primes, containing the primes described in Theorem 8.2, the prime $2$, and the primes so that the conditions of [23, Proposition 5.13] are satisfied for all $p\not\in\mathcal{P}$. Moreover, for all $p\not\in\mathcal{P}$, we define
$\chi_{t}(p):=\left(\frac{-t}{p}\right),\ \psi(p):=\left(\frac{-1}{p}\right).$
If $F$ and $\mathcal{P}_{k,-q,r}$ are as in Theorem 8.2 and $\xi=(1,0,1)^{t}$ (in particular $A(\xi)\neq 0$), we have
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)=A(\xi)L_{\mathcal{P}}\left(F;s-k+3/2\right)\zeta_{\mathcal{P}}(2s-2k+4)^{-1}\frac{\zeta_{\mathcal{P}}(s-k+2,\chi_{t})}{\zeta_{\mathcal{P}}(s-k+2,\psi)}.$
@@
The proof follows by Theorem 8.2 after choosing $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. We have computed $\zeta_{\xi,\mathcal{P}}(s)$ in (20) and
@@
as this can be computed by [23, Proposition 5.15], because $\mathcal{P}$ is chosen so that conditions of [23, Proposition 5.13] are satisfied. ∎
@@
In the case $t=1$, we recover (partly) the result of Kohnen and Skoruppa in [13]. In particular, it is clear that with the above approach, some Euler factors might be missing. However, the benefit is that we also obtain results for $t>1$. These could be interpreted as results for modular forms on a paramodular group (cf. [15, Corollary 6], [4]).
@@ 9.2. The rank $n\geq 2$ case
In the rank $n\geq 2$ case, a Theorem like [26, Theorem 5.7] is not available. For this reason, we seek examples of matrices $S$ so that the following conditions hold:
@@
    The number of classes in the genus of maximal lattices (this is independent of the choice of the maximal lattice, see [24, Paragraph 9.7]) in $H(\xi)_{\mathbb{Q}}=\hbox{SO}^{\psi}(W)$ is $1$. Here, $\psi:=\phi_{0}\mid_{W}$.
We will show that for rank $n=2,4,6,8$, there is at least one positive definite even symmetric matrix $S$ of rank $n$, so that the above conditions are satisfied. We start with the following lemma:
@@
We have $H(\xi)_{\mathbb{Q}}=\textup{SO}^{\psi}(W)$ and $\psi$ can be represented by the matrix
$T=\frac{1}{2}\begin{pmatrix}-2&\\ &-S\end{pmatrix}$. (21)
@@
We have $W=\{x\in V_{0}\mid\phi_{0}(x,\xi)=0\}$. Now $\xi\in U:=\mathbb{Q}e_{1}+\mathbb{Q}e_{n+2}$ and $W=(W\cap U)\oplus U^{\perp}$. But on $U^{\perp}$, $\phi$ is represented by $-S$. Moreover, $W\cap U$ has dimension $1$ and if we write $x=\lambda e_{1}+\mu e_{n+2}\in W\cap U$, we have $\phi_{0}(x,\xi)=\lambda+\mu$ and so $W\cap U$ is spanned by $e_{1}-e_{n+2}$. By evaluating $\phi_{0}[e_{1}-e_{n+2}]=-1$, the lemma follows. ∎
@@
  • $n=6$: $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$ with determinant $3$.
  • $n=8$: $S=\begin{pmatrix}2&-1&1&1&-1&-1&1&-1\\ -1&2&0&-1&0&1&-1&1\\ 1&0&2&1&-1&0&0&0\\ 1&-1&1&2&-1&-1&1&-1\\ -1&0&-1&-1&2&1&-1&1\\ -1&1&0&-1&1&2&-1&1\\ 1&-1&0&1&-1&-1&2&-1\\ -1&1&0&-1&1&1&-1&2\end{pmatrix}$ with determinant $1$.
@@
The matrices $S$ of rank $2$ correspond to the unitary groups of the imaginary quadratic fields $\mathbb{Q}(\sqrt{-3}),\mathbb{Q}(\sqrt{-15})$ respectively (see [21, Example 1.6.6, (v)]).
Let us first check the conditions $(1)$ and $(3)$, right before Lemma 9.1. Condition $(1)$ follows by [21, Proposition 1.6.12] for the matrices with square-free determinant and [21, Lemma 1.6.5, (ii)] for the matrix of determinant $9$.
Condition $(3)$ follows by [10, Section 8], because for the above choices of $S$, the matrix $T$ of (21) corresponds to the following quadratic forms:
  • $n=2$: Examples 4, 26 in matrices of $3$ variables in [10, Section 8].
  • $n=4$: Examples $3,5$ in matrices of $5$ variables in [10, Section 8].
  • $n=6$: Example $3$ in matrices of $7$ variables in [10, Section 8].
  • $n=8$: Example $1$ in matrices of $9$ variables in [10, Section 8].
All these examples correspond to one class in the genus of the standard lattice $\mathbb{Z}^{n+1}$, in the cases when it is maximal. This can also be seen by a simple computation in MAGMA for example. As we mentioned above, this shows that every maximal lattice has one class in its genus. Finally, we need to check condition $(2)$. We will use [24, Proposition 11.12]. We have the following local result.
@@
Our proof is based on [24, Proposition 11.12]. We will first establish the following claim:
@@
This follows from the fact that $L_{0,p}^{*}=S_{0}^{-1}L_{0,p}$ and that
$S_{0}^{-1}=\begin{pmatrix}&&1\\ &-S^{-1}&\\ 1&&&\end{pmatrix},\textup{ }S^{-1}=\frac{1}{\det(S)}\textup{adj}(S).$
We remind ourselves that $\phi_{0}[\xi]=1$ and $\phi_{0}(\xi,L_{0,p})=\frac{1}{2}\mathbb{Z}_{p}$ for all primes $p$. Therefore, $\phi_{0}(\xi,L_{0,p})^{2}=\phi_{0}[\xi]\mathbb{Z}_{p}$ for all $p\neq 2$. Moreover, $L_{0,2}^{*}=L_{0,2}$ for every choice of $S$, as $2\nmid\det(S)$ for any $S$. This means [24, Proposition 11.12] is applicable in every case.
Let now $t_{p}$ denote the dimension of the maximal anisotropic subspace of $(\mathbb{Q}_{p}^{n+2},\phi_{0})$ (cf. [24, Paragraph 8.3]). For $p\nmid\det(S)$, we have $L_{0,p}^{*}=L_{0,p}$, $t_{p}\neq 1$ as $n$ is even and $4\phi_{0}[\xi]^{-1}\phi_{0}(\xi,L_{0,p})^{2}=\mathbb{Z}_{p}$. So, by [24, Proposition 11.12, (iii), (2)], $L_{0,p}\cap W_{p}$ is $\mathbb{Z}_{p}$-maximal.
If now $p\mid\det(S)$, we claim $t_{p}>1$ and therefore [24, Proposition 11.12, (iii), (1)] will be applicable. We show this case by case. Define $K_{0}:=\mathbb{Q}_{p}(\sqrt{\delta})$, where $\delta:=(-1)^{(n+2)(n+1)/2}\det(\phi_{0})$. We note that from the proof of [25, Lemma 3.3], we have $t_{p}=2$ if and only if $K_{0}\neq\mathbb{Q}_{p}$.
  • $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. Then, we claim that $t_{3}=2$. Now, $\det(\phi_{0})=-3/2^{4}$ and then $K_{0}=\mathbb{Q}_{3}(\sqrt{\det(\phi_{0})})=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$.
  • $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, $\det(\phi_{0})=-15/2^{4}$ and we claim $t_{3}=t_{5}=2$. But again, if $p\in\{3,5\}$, we have $K_{0}=\mathbb{Q}_{p}(\sqrt{\det{\phi_{0}}})=\mathbb{Q}_{p}(\sqrt{-15})\neq\mathbb{Q}_{p}$ by taking valuations (for example, if $\sqrt{-15}\in\mathbb{Q}_{3}$, then $-15=u^{2}$ for some $u\in\mathbb{Q}_{3}$ and so $2v_{3}(u)=1$, contradiction).
  • $S=\begin{pmatrix}2&-1&-1&-1\\ -1&2&1&0\\ -1&1&2&0\\ -1&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-5/2^{6}$ and so $K_{0}=\mathbb{Q}_{5}(\sqrt{5})\neq\mathbb{Q}_{5}$, so $t_{5}=2$.
  • $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, $K_{0}=\mathbb{Q}_{3}$, and using a software (e.g. SageMath) we can compute $t_{3}=4$.
  • $S=\begin{pmatrix}2&1&-1&1&-1&1\\ 1&2&0&1&-1&1\\ -1&0&2&-1&1&0\\ 1&1&-1&2&-1&0\\ -1&-1&1&-1&2&0\\ 1&1&0&0&0&2\end{pmatrix}$. In this case, $\det(\phi_{0})=-3/2^{8}$ and so again $K_{0}=\mathbb{Q}_{3}(\sqrt{-3})\neq\mathbb{Q}_{3}$. Therefore, $t_{3}=2$.
@@
Finally, the fact that $D_{p}=H(\xi)_{p}\cap C_{p}$ follows from [24, Proposition 11.12, (iv)], as $t_{p}\neq 1$ always, because $n$ is even. ∎
@@
The first claim follows by [21, Proposition 1.6.9], as all the localizations are maximal by Proposition 9.2. For the second one, we have
@@
But the lattice $xL_{0}$ is the lattice which is defined by $(xL_{0})_{p}=x_{p}L_{0,p}$ for all primes $p$. Now, if $x\in H(\xi)_{\mathbb{A}}$ with $xL_{0}=L_{0}$, we have $x(L_{0}\cap W)=L_{0}\cap W$ (see [24, page 104]), so $H(\xi)_{\mathbb{A}}\cap C\subset D$.
On the other hand, for all primes $p$, let $D_{p}:=D\cap H(\xi)_{p}$. Then, if $x\in D$, then $x_{p}\in D_{p}$, so $x_{p}\in H(\xi)_{p}\cap C_{p}$ by Proposition 9.2. Therefore, $x_{p}L_{0,p}=L_{0,p}$ for all primes $p$. This means that $x\in H(\xi)_{\mathbb{A}}\cap C$, as wanted. ∎